# Bilevel subsidy-enabled mobility hub network design with perturbed utility coalitional choice-based assignment

### v1

@@ 3.4 Solution method
6:  Step 3: Load and warm start the solution process for Eq. (16) using $\mathbf{x^{k},v^{k},\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ and $\rho^{k}$. Update the bounds for $\mathbf{\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ by $\zeta$. Solve until time limit $T$ is reached or the optimality gap is lower than $\tau$. Obtain the new solution $\mathbf{p}$.
@@
19:  Step 6: $\textbf{p}^{k}=\textbf{p}$
@@
22:Solve lower level problem Eq. (1) with $\textbf{p}^{*}$. Obtain $\mathbf{x}^{*},\mathbf{v}$
23:Output: Optimal solution $(\mathbf{p}^{*},\mathbf{x}^{*},\mathbf{v}^{*})$.
@@ 4 Numerical experiments
In this section, we provide two sets of experiments to illustrate the formulation and the solution algorithm. We first use a toy network to verify the algorithm and illustrate the model. The second result is tested on a hypothetical multimodal network centered around three LIRR stations, which serve as potential mobility hubs for travelers commuting to New York City from Long Island.
@@ 4.1 Illustrative example
A toy network is illustrated in Fig. 2. Three OD pairs are considered: (1, 0), (2, 0), and (3, 0), each with a demand of 100 units over a typical operating period. The MaaS platform is considered for the transit line from station S to node 0, as well as MOD service links provided by a single operator from nodes 1, 2, and 3 to node H to connect to station S. The gray dotted links represents non-MaaS participating services (e.g., privately owned vehicles, walking, other non-MaaS services). For this illustrative case, only the MH node $H$ has the capacity constraint. The MH capacity reflects the supply side resources for supporting a fleet there: dedicated space for temporary stationing, maintenance, fueling, etc.
@@
All input parameters are listed in Table 2. All access and egress links are assigned with a small length value, which is used to prevent infeasible solutions without impacting the flow assignment results. The three MOD service links (A-H, B-H, C-H) are controlled by the MaaS platform to adjust suitable service pricing in the upper level problem, while other services having fixed pricing aggregated with the the travel cost. $c_{l}^{t},c_{l}^{o},p_{l}$ are all measured in monetary values. The lower level problem decides the joint choice of traveler flow and hub capacities.
@@
The toy example is solved on a device with an Intel(R) i7-13705 processor. We directly solve it to global optimality using Gurobi 12.1 and the branch-and-bound algorithm, with the optimality gap being less than $10^{-6}$. All scenarios reaches global optimum under 0.5 seconds.
BENCHMARK SCENARIO: To better understand how the price allocation scheme impact both the MH based service flow and platform revenue, we run a benchmark scenario representing a traditional transit hub where operators may converge for transfers, but no digitalization is setup for cost transfers between the operators. This solution is obtained by simply solving the lower-level model only with a fixed price of \$3\\$3. When MOD service link prices are all fixed to \$3 per unit distance, only link (A, H) is actively used with a flow of 24.47 from node 1. Both nodes 2 and 3 are not served by the MOD service. Total revenue is \$73.4 and the proportion of opened MH capacity is 12.24%.
BASE SCENARIO: The results are summarized in Table 3 under the columns ”Price 1” and ”Flow 1.” OD pair (3, 0) exclusively relies on the direct non-MaaS link, whereas OD pairs (1, 0) and (2, 0) exhibit multimodal routing behavior that combines both direct and MaaS-integrated links. Notably, OD (1, 0) distributes its flow across three routes: the direct link, a non-MaaS path using the transit connection (1S0)(1-S-0), and a MOD-only path via (1AHS0)(1-A-H-S-0), highlighting the benefits to travelers from node 1 when MOD services are introduced. The MaaS-based MOD service links (A, H) and (B, H) are priced at 3.0 and 0.937, respectively, to maximize total revenue. The latter reflects a subsidy of \$2.06 per unit distance for travelers using MOD service from zone 2. For OD pair (3, 0), the direct link in the dummy subnetwork dominates all other options. even when the platform offers free access to the MOD service link (C, H), no travelers would choose to use it nor would the operator serve that area. Due to the overall low usage of MaaS services in this setting, the hub is heavily underutilized, with a capacity set to only 16.9% of the maximum capacity. The total MaaS service revenue in this case is \$90.8, a 23.7% increase over the benchmark. The value of having a cyberphysical MH with enabled subsidy is equal to \$16.6 per operating period.
Remark 1.
Compared to a benchmark transit hub, the subsidy-enabled MH can produce a higher capacity and increased revenues; the value of having a such a MH can be quantified as the difference in the objective value.
ALTERNATIVE SCENARIO: Now consider a scenario where the service price is capped to \$2\\$2 per unit distance. As shown in the ”Price 2” and ”Flow 2” columns of Table 3, more travelers use the MOD service links on link (A, H) due to the lowered cap. For link (B, h), the price level remains the same, although with the price cap the required subsidy is reduced to only \$1.06. Again, no travelers choose link (C, H) due to the dominance of the direct link (3, 0). As a result, the platform sets the capacity to 20.31% of maximum hub capacity to serve MaaS users. The MaaS platform earns a total of \$80. The \$10 dollar loss in revenue due to further constraining the maximum price translates to a higher enrollment in users coming from zone 1 at an overall lower social surplus.
Remark 2.
The model captures the social surplus and ridership impacts of a price cap.
@@
Source node Sink node $\mathbf{c_{l}^{t}(\$)}$ $\mathbf{c_{l}^{o}(\$)}$ $\mathbf{\hat{p}_{l}(\$)}$ $\mathbf{d_{l}}$
1 0 7 0 N/A 4
2 0 7 0 N/A 4
3 0 7 0 N/A 4
A H 4 1 3 1
B H 4 1 3 2
C H 4 1 3 3
1 S 6 0 N/A 1
2 S 6 0 N/A 2
3 S 6 0 N/A 3
H S 0 0 N/A 0.1
S 0 6 0 N/A 4
1 A 0 0 N/A 0.1
2 B 0 0 N/A 0.1
3 C 0 0 N/A 0.1
Maximum hub capacity, $\mathbf{z_{i}}$ 200
$\mathbf{\alpha}$ 0.5
Capacity cost, $\mathbf{c_{i}}$ 1
Table 3: Comparison of link prices and flows across base and alternative scenarios, with subsidies in ( )
Link OD Pair $s$ Price 1, $p_{l}$ ($r_{l}$) Flow 1, $x_{l,s}q_{s}$ Price 2, $p_{l}$ ($r_{l}$) Flow 2, $x_{l,s}q_{s}$
(1, 0) (1, 0) N/A 53.07 N/A 49.28
(2, 0) (2, 0) N/A 90.71 N/A 90.71
(3, 0) (3, 0) N/A 100 N/A 100
(1, A) (1, 0) N/A 24.47 N/A 31.30
(A, H) (1, 0) 3 24.47 2 (0) 31.30
(H, S) (1, 0) N/A 24.47 N/A 31.30
(2, B) (2, 0) N/A 9.29 N/A 9.29
(B, H) (2, 0) 0.94 (2.06) 9.29 0.94 (1.06) 9.29
(H, S) (2, 0) N/A 9.29 N/A 9.29
(1, S) (1, 0) N/A 22.46 N/A 19.42
(S, 0) (1, 0) N/A 46.93 N/A 50.72
(S, 0) (2, 0) N/A 9.29 N/A 9.29
Hub Capacity, $v_{i}$ 16.88% 20.31%
Objective value, $\Phi_{0}$ \$90.8\\$90.8 \$80\\$80
@@ 4.2 LIRR mobility hubs case study
We further use three LIRR stations: Ronkonkoma, St. James, and Sayville, and their surrounding neighborhoods to test the proposed model. These three stations are treated as candidate mobility hubs. We use the centroids of census tracts within a 5-mile radius of each station as origin nodes, with Manhattan as the destination for a typical weekday morning AM period. Two layers of subnetworks are defined, similar to the toy network. Fig. 3 illustrates the MaaS subnetwork. Each service node is directly connected via MOD service links, with candidate mobility hubs established at each LIRR station. In total, 78 service nodes are included in the network with 113 direct microtransit service links. Additionally, three LIRR links are included to represent line-haul services.
@@
The MOD service considered in this case study is a hypothetical new microtransit operation that operates at unit distance fare rate of \$2\\$2 per mile. For the dummy subnetwork, 78 direct service links parallel to the MOD links are added to represent other access services. Another 78 direct links from service nodes to the Manhattan node are also included in the dummy subnetwork. As a result, the whole network has 244 nodes and 469 links serving the 78 OD pairs. By comparison, the expanded Sioux Falls network in Liu and Chow, (2024) has 30 OD pairs, 82 nodes, and 748 links.
@@
The 78 OD pairs, each originating from a service node to Manhattan, have simulated demand drawn from a normal distribution with a mean of 60 and a standard deviation of 20, representing a typical weekday AM period. The simulated total demand is 4,734, with an average being 60.69 per census tract OD pair. Note that this level of demand falls in the range of observation, as Ronkonkoma Station, the largest one, has 5452 free parking spaces for general passengers, and fills up on a typical weekday. Other input elements are summarized in Table 4. $\bar{d}_{l}$ represents the average link length.
For experimental design, we first use a base case that optimizes link price in the upper level to determine the objective value. The following tests are conducted. (1) To quantify the value of a MH at Ronkonkoma, we then remove the Ronkonkoma station and corresponding MOD links and compare the resulting optimal objective value with the base case. (2) Afterwards, we evaluate a scenario where the microtransit provider increases their price from \$2 to \$3 per mile. (3) The platform considers an alternative objective of maximizing ridership (passenger-miles) under a budget subsidy of \$5000. In this case, we use Eq. (14) to solve problem. (4) Finally, the platform also considers setting the subsidy by MH instead of by link to observe the changes in flow.
@@
All cases are run on a device with an Intel(R) i7-13705 processor. We initiate the proposed algorithm with $\rho_{0}$ equals to 500 and terminate the algorithm either finishing 10 iterations or the optimality gap is below 0.1%. Each iteration has a runtime limit of 10 minutes.
@@
Subnetwork Mode $\mathbf{c^{t}_{l}}$ $\mathbf{c^{o}_{l}}$ $\mathbf{\hat{p}_{l}}$ $\mathbf{\bar{d}_{l}}$
MOD MOD 1.00 0.5 2 5
LIRR 0.80 0 N/A 50
Dummy LIRR dummy 0.90 0 N/A 50
Drive 0.95 0 N/A 60
Park-n-Ride 1.20 0 N/A 5
@@ 4.2.1 Base case
The proposed algorithm finds a solution with an optimality gap of less than 0.1% in under 300 seconds. The runtime is significantly shorter than other link-based MaaS design models proposed by previous studies (Liu and Chow,, 2024; Yao and Zhang,, 2024) when solving problems with similar scales. In comparison, the Sioux-Fall network-based case described in Liu and Chow, (2024) requires longer than a 4-hour run time to converge. This illustrates the potential of the proposed algorithm in real-world applications.
The final solution has all MOD service links priced at \$2 per unit length, which is the upper limit of the service price (Remark 3). The total revenue from the MOD service is \$15,959, and 34.19% (1619 passengers) of the total OD demand uses the microtransit service. 752 people use service base at Ronkonkoma station, 583 use service base at St.James station, and 401 use service base at Seyville. This solution suggests a MH platform design with subsidy provision is unnecessary as the pricing offered by the existing microtransit operator is satisfactory to travelers. In other words, there is sufficient incentive for the microtransit operator to enter this last mile market by simply providing service in these areas without further incentives from MaaS fare bundling.
@@
The model can inform policymakers on the need for a subsidy-enabled MH platform design or to simply implement a traditional transit hub.

### v2

@@ 3.4 Solution method
6:  Step 3: Load and warm start the solution process for Eq. (15) using $\mathbf{x^{k},v^{k},\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ and $\rho^{k}$. Update the bounds for $\mathbf{\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ by $\zeta$. Solve until time limit $T$ is reached or the optimality gap is lower than $\tau$. Obtain the new solution $\mathbf{p},\mathbf{r}$.
@@
19:  Step 6: $\textbf{p}^{k},\textbf{r}^{k}=\textbf{p},\textbf{r}$
@@
22:Solve lower level problem Eq. (1) with $\textbf{p}^{*},\textbf{r}^{*}$. Obtain $\mathbf{x}^{*},\mathbf{v}^{*}$
23:Output: Optimal solution $(\mathbf{p}^{*},\mathbf{r}^{*},\mathbf{x}^{*},\mathbf{v}^{*})$.
@@ 4 Numerical experiments
In this section, we provide two sets of experiments to illustrate the formulation and the solution algorithm. We first use a toy network to verify the algorithm and illustrate the model. The second experiment is tested on a hypothetical multimodal network centered around three LIRR stations, which serve as potential mobility hubs for travelers commuting to New York City from Long Island.
@@ 4.1 Illustrative example
A toy network is illustrated in Fig. 2.Three OD pairs are considered: (1, 0), (2, 0), and (3, 0), each with a demand of 100 units over a typical operating period. The MaaS platform is considered for the transit line from mobility hub H to node $0^{\prime}$, as well as MOD service links from nodes 1, 2, and 3 to node H. For node 1, two MOD service options are provided (e.g., car sharing and micromobility) with different cost attributes for both travelers and operators. The gray dotted links represent out-of-platform services (e.g., privately owned vehicles, walking, or other pay-as-you-go services). All platform-based nodes A, B, C, D, and H have virtual links to represent node capacities. Nodes A, B, and C are operated by the car sharing company (denoted as MOD 1) and node D is operated by the micromobility company (denoted as MOD 2). Both the MH node and MOD service nodes have capacity constraints and associated capacity costs. The MH capacity reflects the supply-side resources for supporting a fleet there: dedicated space for temporary stationing, maintenance, fueling, etc. The service node capacity reflects the availability constraint and the associated service stop establishment cost. For each OD pair, dedicated platform access links are defined for traveler pricing purposes.
@@
Core input parameters are listed in Table 2. All access, egress, and virtual capacity links are assigned a small length value, which is used to prevent infeasible solutions without impacting the flow assignment results. Virtual node capacity links have zero cost and price and are also assigned a small length value to prevent infeasibility. The three MOD access links ($1-1^{\prime}$, $2-2^{\prime}$, $3-3^{\prime}$) are controlled by the platform to adjust suitable service pricing in the upper-level problem along with operator-facing subsidies assuring operator profitability. Other services have fixed costs. $c_{l}^{t}$, $c_{l}^{o}$, and $p_{l}$ are all measured in monetary values. We want to highlight that after paying the access fee when traversing link $1-1^{\prime}$, travelers from node 1 are free to choose either $1^{\prime}-A$ or $1^{\prime}-D$ without additional fare cost, which reflects the ”bundled” feature of such a platform. Such behavioral decisions are determined by the lower-level problem, which decides the joint choice of traveler flow and hub capacities.
@@
The toy example is solved on a device with an Intel(R) i7-13705 processor. We directly solve it to global optimality using Gurobi 12.1 and the branch-and-bound algorithm, with the optimality gap being less than $10^{-4}$. All scenarios reach the global optimum under 1 second.
The results are summarized in Table 3 under the columns ”Price”, ”Subsidy”, and ”Flow”. Flows on virtual capacity links (e.g., D-D’) are not shown in the table for simplicity. OD pair (3, 0) relies exclusively on the direct non-MaaS link, whereas OD pairs (1, 0) and (2, 0) exhibit multimodal routing behavior, combining both direct and MaaS-integrated links. Specifically, flows for OD (1, 0) and (2, 0) are distributed across three routes: the direct link, a non-MaaS path using the transit connection, and a MOD-only path. This highlights the benefits to travelers starting from nodes 1 and 2 when MOD services are introduced. However, under the provided pricing scheme for OD (1, 0), no travelers select service node A, and the service link chain $A-A^{\prime}-H$ remains closed. This reflects the competitive dynamics of the market, where multiple services may be available but not all are viable for platform operation.
The platform access fees for OD (1, 0) and (2, 0) are set at 8.49 and 6.84, respectively, to maximize total revenue. To incentivize operators to participate in the platform without incurring losses, the operator receives a subsidy of 3.50 per traveler entering the service at node D for OD (1, 0). For OD pair (2, 0), the corresponding operator subsidy is 4.50. This indicates that involving operators in the route from service node B to hub H is more costly for the platform. For nodes A and C, other modes dominate the options. Therefore, no subsidy is offered, as opening services at these nodes is deemed unprofitable.
For OD (3, 0), the direct link in the dummy subnetwork dominates all other options. Even if the platform were to offer free access to the MOD service link (C,H), travelers would not choose it, nor would the operator serve that area. Due to the overall low usage of MaaS services in this scenario, the hub is significantly underutilized, operating at only 21.59% of its maximum capacity. The total platform service profit in this case is \$161.10.
@@
Source Sink $\mathbf{c_{l}^{t}(\mathdollar)}$ $\mathbf{c_{l}^{o}(\mathdollar)}$ $\mathbf{\hat{p}_{l}(\mathdollar)}$ $\mathbf{d_{l}}$
1 0 9 0 N/A 4
2 0 9 0 N/A 4
3 0 9 0 N/A 4
1 $1^{\prime}$ 0 0 10 0.1
2 $2^{\prime}$ 0 0 10 0.1
3 $3^{\prime}$ 0 0 10 0.1
$1^{\prime}$ A 0 0 N/A 0.1
$1^{\prime}$ D 0 0 N/A 0.1
$2^{\prime}$ B 0 0 N/A 0.1
$3^{\prime}$ C 0 0 N/A 0.1
$A^{\prime}$ H 5 2 N/A 1
$B^{\prime}$ H 5 2 N/A 2
$C^{\prime}$ H 5 2 N/A 3
$D^{\prime}$ H 3 3 N/A 1
1 S 7 0 N/A 1
2 S 7 0 N/A 2
3 S 7 0 N/A 3
$H^{\prime}$ $0^{\prime}$ 6 0 N/A 4
S 0 7 0 N/A 4
Maximum hub capacity, $b_{l}$ 200
Maximum service node capacity, $z_{l}$ 50
Maximum subsidy allowed per traveler, $\hat{r}_{l}$ 5
$\mathbf{\alpha_{1}},\mathbf{\alpha_{2}}$ 1, 0.5
Capacity cost, $c_{i}$ 1
Table 3: Comparison of link prices, subsidies and flows across base and alternative scenarios
Link OD Pair $s$ Price, $p_{l}$ Subsidy, $r_{l}$ Flow, $x_{l,s}q_{s}$
(1, 0) (1, 0) N/A N/A 32.87
(2, 0) (2, 0) N/A N/A 82.59
(3, 0) (3, 0) N/A N/A 100
(1, 1’) (1, 0) 8.49 N/A 30.82
(1’, D) (1, 0) N/A 3.5 30.82
(D’, H) (1, 0) N/A N/A 30.82
(H’, 0) (1, 0) N/A N/A 30.82
(2, 2’) (2, 0) 6.84 N/A 12.35
(2, B) (2, 0) N/A 4.5 12.35
(B’, H) (2, 0) N/A N/A 12.35
(H’, 0) (2, 0) N/A N/A 12.35
(1, S) (1, 0) N/A N/A 36.30
(S, 0) (1, 0) N/A N/A 36.30
(2, S) (2, 0) N/A N/A 5.06
(S, 0) (2, 0) N/A N/A 5.06
Hub Capacity, $v_{i}$ 21.59%
Objective value, $\Phi_{0}$ $\mathdollar 161.10$
@@ 4.2 LIRR mobility hubs case study
We further use three LIRR stations: Ronkonkoma, St. James, and Sayville, and their surrounding neighborhoods, represented by travel analysis zones (TAZs), to test the proposed model. These three stations are treated as candidate mobility hubs. We use the centroids of census tracts within a 5-mile radius of each station as origin nodes, with Manhattan as the destination for a typical weekday morning AM period. In total, 78 neighborhoods are covered. Two layers of subnetworks are defined, similar to the toy network. Fig. 3 illustrates the platform subnetwork. Each service node is directly connected via MOD service links, with candidate mobility hubs established at each LIRR station. We assume these primary MOD services are managed by a single operator. To model market competition similar to the toy network, we introduce a second MOD service provider in the vicinity of Ronkonkoma station. This secondary operator functions within a 3-mile radius, covering 20 neighborhoods. In a real world context, the primary MOD service (MOD 1) represents larger scale microtransit options, such as ride hailing or ridepooling. For example, Suffolk County operates an on-demand service in Southampton that provides last mile access to the LIRR Southampton station and the S92 bus (Southampton, 2026). The localized MOD service (MOD 2) represents micromobility solutions, such as bike sharing or shared scooters. The resulting network comprises 98 service nodes connected by 133 direct microtransit service links. Additionally, three LIRR links are included to represent line-haul services.
@@
For the dummy subnetwork, 78 direct service links parallel to the MOD links are added to represent other access services. Another 78 direct links from service nodes to the Manhattan node are also included in the dummy subnetwork. As a result, the whole network has 380 nodes and 642 links serving the 78 OD pairs. By comparison, the expanded Sioux Falls network in Liu and Chow (2024) has 30 OD pairs, 82 nodes, and 748 links.
@@
The 78 OD pairs, each originating from a service node to Manhattan, have demand simulated from a normal distribution with a mean of 60 and a standard deviation of 20, representing a typical weekday AM period (see Table. LABEL:tab:simulated_demand in Appendix for details). The simulated total demand is 4,734, with an average being 60.69 per census tract OD pair. Note that this level of demand falls in the range of observation, as Ronkonkoma Station, the largest one, has 5,452 free parking spaces for general passengers, and fills up on a typical weekday.
Each MOD service node is assigned a capacity of 50, while the Mobility Hub (MH) has a capacity of 2,000. The maximum platform access fee (which includes the LIRR ride) is capped at \$30, and the operator subsidy is limited to a maximum of \$5 per traveler. $\alpha_{1}$ and $\alpha_{2}$ are set to be 1 and 0.5, respectively. Other input elements are summarized in Table 4. The interpretation of the $\hat{p_{l}}$ parameter varies depending on the link type. For MOD platform access links, it represents the price cap. Conversely, for dummy network links, it represents the fixed fee paid by travelers. $\bar{d}_{l}$ represents the average link length. To reflect the trade-off between micromobility and microtransit, MOD 2 is modeled with lower operating costs but slower service speeds than MOD 1.
For the experimental design, we first establish two baseline scenarios: one without the proposed platform (”business as usual”) by solving the lower level problem without the platform subnetwork, and a second with the proposed platform by solving the problem with the stated network and input variables to determine the baseline objective value. We then perform a sensitivity analysis focusing on MOD operating costs within the platform network and driving costs within the dummy network to illustrate the sensitivities captured by the model. To further evaluate system dynamics, we then analyze three targeted scenarios: (1) eliminating micromobility links to assess operator participation effects; (2) removing the Ronkonkoma Mobility Hub designation to quantify its added value; and (3) reducing the operator subsidy cap to \$3 to examine the impact of tighter fiscal constraints on planning decisions.
@@
All cases are run on a device with an Intel(R) i7-13705 processor. We initiate the proposed algorithm with $\rho_{0}$ equal to 300 and terminate the algorithm either after 10 iterations or when the optimality gap is below 0.1%. Each iteration has a runtime limit of 10 minutes.
@@
[…]
MOD Access 0 0 20 0.1
MOD 1 0.42 0.5 0 3
MOD 2 0.5 0.3 0 1.8
LIRR 1.2 0 0 25
Dummy LIRR dummy 1.30 0 20 25
Drive 1.33 0 21 30
Park-n-Ride 0.5 0 0 5
@@ 4.2.1 Base case
Prior to solving the full bilevel model, we solved the lower-level model exclusively on the dummy network to establish traveler behavior in the absence of platform participation. Using the parameters listed in Table 4, we observe a mode split where 64.52% of travelers utilize the existing transit network and 35.45% opt for direct driving. The lower level objective value is 5,451.97. These results serve as a ”business as usual” baseline for comparison with the proposed bilevel model, allowing us to quantify the impact of introducing the platform service.
@@
We then run the full bilevel model with the platform network added. The proposed algorithm finds a solution with an optimality gap of less than 1% in $3092$ seconds. The runtime is significantly shorter than other link-based MaaS design models proposed by previous studies (Liu and Chow, 2024; Yao and Zhang, 2024) when solving problems with similar scale. In comparison, the Sioux-Fall network-based case described in Liu and Chow (2024) requires longer than a 4-hour run time to converge. This illustrates the scalability of the proposed algorithm in real-world applications.
To demonstrate convergence to the global optimum, we solved the model across a range of $\rho$ values. We then fix the upper-level decision variables to their corresponding optimal outputs and solved the lower-level model directly. Using the resulting flows and capacities, we calculated the actual platform profit. As illustrated in Fig. 4, both values converge as $\rho$ increases. However, the rate of convergence diminishes as $\rho$ approaches a critical value. Once $\rho$ exceeds this threshold, the optimum of the single-level formulation effectively represents the true global optimum of the bilevel model. Future studies could focus on developing algorithms that efficiently identify this critical threshold to further accelerate the solution process.
Figure 4: Convergence between single level objective and true platform profit.
The final solution has platform access prices (which includes the access mode and LIRR) ranging from \$22.47 to \$30, and the general distribution can be observed in Fig. 6(a). This pricing structure is economically reasonable when compared to general LIRR commuting costs, ensuring the platform remains a viable option for travelers. The total profit from the MOD service is \$38721.23, and 35.32% (1,661 passengers) of the total OD demand uses the MH centered platform. 741 people use service based at Ronkonkoma station, 550 use service based at St. James station, and 380 use service based at Sayville. Fig. 5 illustrates the percent mode share of the MOD services across the 78 neighborhoods.
Figure 5: Mode share of MH based MOD services.
The operator subsidy ranges from \$0 to \$5. However, the distribution is polarized, with most values falling either below \$1 or above \$4, as illustrated in Fig. 6(b).This indicates that the platform strategically concentrates financial incentives on specific, high-priority service links rather than distributing them uniformly. We observe a strong positive correlation of 0.66 between operator subsidies and passenger flow. This suggests that the platform allocates higher per-traveler subsidies to high-demand links to ensure that the total subsidy volume is sufficient to cover the operator’s aggregate costs. Conversely, the correlation between link length and subsidy level is weakly positive, at 0.11. While areas further from MHs generally require higher subsidies to be viable, distance alone is not the sole determinant. Areas easily accessible via the dummy network often have lower platform usage. Furthermore, variations in access fees create different passenger flows for areas at similar distances. Consequently, the relationship between distance and subsidy is only weakly correlated.
Figure 6: Histogram of (a) platform access fee and (b) histogram of operator subsidy.
Both the major operator (MOD 1 in Table 2) and the smaller operator receive the subsidies to reach break even. The major operator incurs \$4781.41 in operating costs for the in-platform service and receives an equivalent \$4781.41 in subsidies, and the smaller operator incurs \$220.76 in operating costs and receives \$220.76 from the platform. With the introduction of the platform service, the mode share of direct driving drops to 22.46%, representing a 36.6% decrease compared to the no-platform scenario. To quantify the utility impact, we compared the lower-level objective values of both models. The objective value decreases from 5,451.97 in the baseline (dummy network only) scenario to 5,154.90 in the optimal platform case. Since the lower-level objective measures user disutility, this reduction demonstrates the platform’s positive impact on overall system welfare. Per Remark 1, this particular baseline suggests there is an opportunity for a subsidy-enabled MH platform to serve this market.
@@
The model identifies an opportunity for a MH platform when it reduces the overall travel disutilities with a fare bundle that can cover the costs of all participating operators.
We subsequently conducted two sensitivity analyses focusing on cost-related components: the MOD link operating cost and the direct driving cost. These analyses reveal how fluctuations in these costs impact both the platform’s profitability and the utilization of the three MHs.
MOD link operating cost sensitivity analysis
We first perform a sensitivity analysis on the operating costs of MOD links. We varied the operating costs for services operated by both MOD 1 and MOD 2 across a range of -50% to +50%. Fig. 7 illustrates the relationship between these operating costs and the platform profit. As operating costs rise, the platform is required to provide higher subsidies. This leads to an increase in the platform access fee, as shown in Table. 5, which subsequently reduces platform service usage (and vice versa). Notably, we observed a linear relationship between operating costs and platform profit. A 50% fluctuation in operating cost results in a 5% shift in platform profit under these baseline settings. Given that the subsidy comprises only 13% of the total platform revenue in the base scenario, even a 50% variation in cost leads to only a marginal adjustment in the platform access fee.
Figure 7: Relationship between platform profit and MOD link operating cost
Table 5: Impact of operating cost changes on access fee and share of platform usage
Operating Cost Change (%) Average Access Fee Relative Fee Change (%) Share (%) of MH usage
-50 26.13 -1.94 36.14
-25 26.37 -1.06 35.79
0 26.65 0.00 35.32
+25 26.92 1.01 34.82
+50 27.16 1.93 34.30
Fig. 8 illustrates the relationship between link operating costs and the volumes across the three MH locations. While the volumes at all three hubs exhibit a consistent inverse relationship with operating costs, the response is asymmetric. The negative impact of a cost increase is stronger than the positive gain from a cost reduction. Specifically, a 50% decrease in operating cost increases hub volumes by 2% to 2.5%, with St. James benefiting the most. However, a 50% cost surge leads to a sharper decline of 3.0%, with Ronkonkoma becoming the most impacted. This suggests that while lower costs encourage some additional usage, higher costs penalize demand more severely.
Figure 8: Relationship between MH usage and MOD link operating cost
Direct driving cost sensitivity analysis
We subsequently perform a sensitivity analysis on direct driving costs to evaluate the impact of competing mode pricing. Fig. 9 illustrates the resulting relationship. Since private driving functions as a substitute mode, increasing its cost dissuades travelers from using private vehicles. This drives higher demand for MH services and increases platform profit. For instance, a 43% increase in driving costs, which is comparable to a \$9 congestion toll for entering New York City, results in a 16% gain in platform profit. Consistent with our operating cost analysis, the relationship between platform profit and driving cost changes is linear. Table. 6 details the corresponding shifts in access fees and the platform’s overall mode share. Notably, the platform access fee covaries positively with driving costs. As driving becomes more expensive, the platform strategically raises its access fee to capitalize on the influx of deterred drivers. This creates an indirect effect where travel becomes universally more expensive, negatively impacting travelers who rely on the platform as an alternative to driving.
Figure 9: Relationship between platform profit and direct driving cost
Table 6: Impact of driving cost changes on access fee and platform mode share
Driving Cost Change (%) Avg. Access Fee Relative Fee Change (%) Mode Share (%)
-52 24.59 -7.72 31.66
-43 25.00 -6.17 32.27
-33 25.42 -4.63 32.88
-24 25.83 -3.08 33.48
-14 26.19 -1.71 34.17
0 26.65 0.00 35.32
10 26.93 1.06 36.10
19 27.21 2.11 36.89
29 27.49 3.16 37.68
43 27.88 4.60 38.90
Fig. 10 illustrates the relationship between direct driving costs and traffic volumes across the three MH locations. When driving costs decrease, the relative decline in volume is nearly identical across all three hubs. However, as driving costs rise, the relative growth rates begin to diverge slightly. St. James exhibits the greatest relative gain, while Ronkonkoma shows the lowest. Despite this minor divergence, the three hubs demonstrate broadly similar demand elasticity with respect to driving costs. Consequently, policy measures that change the cost of driving would likely produce a consistent impact on ridership across the three hubs.
Figure 10: Relationship between MH usage and direct driving cost
4.2.2 Evaluating the impact of in-platform service competition
To examine the impact of reduced operator competition, we eliminated MOD 2 service entirely. This consolidation resulted in a 0.73% decline in total platform profit (\$38,439.05) and a marginal 0.1% drop in ridership (1,659 passengers). While the average platform access fee remained relatively stable (increasing from \$26.65 to \$26.72), the pricing distribution shifted significantly. The first quartile rose from \$22.95 to \$25.29, and the median increased from \$25.86 to \$27.88. This indicates that reduced competition disproportionately raised costs for lower-priced trips.
Due to the higher operating costs associated with the remaining operator, the total subsidy rises from \$5002.17 to \$5072.32, an increase of 1.40%. Therefore, incorporating more service options into the platform proves beneficial for both the platform and travelers. The platform can potentially reduce total subsidy costs, while travelers benefit from greater access and a wider range of choices within the platform ecosystem with lower access fee.
Remark 2.
Reduced competition in participating access operators to the MH platform can disproportionately raise costs for travelers even as the impact to the platform’s profits may be minimal.